# Native Design Bias: Studying the Impact of English Nativeness on Language Model Performance

### v1

@@ 5.2 Generative tasks
(a) Native vs. non-native
(b) Western native vs. not
Figure 5: Overall average BLEU score of the generation tasks per model and group.
On average, we do not see a clear performance difference for all generative tasks between the native English speakers and the non-native English speakers. Figure 5 shows a slightly higher performance for Qwen for non-native English speakers and both Qwen and Haiku perform better for Western native English speakers. All figures display only the BLEU score, but similar trends are shown by the other metrics. All performance metrics are included in Appendix L.
When analyzing these results per dataset, we find that the best-performing group depends on the specific task at hand. In general, the generative tasks can be divided into two groups: group one with CODA19, which includes a medical research article generation task, and the paraphrasing dataset, group two with StoryCloze and CNN Dailymail. We find that all models prefer the non-native English speakers or not Western native English speakers for group one. Most models prefer (Western) native English speakers for group two. Detailed results are included in Table 17 in Appendix.
In interpreting these results, it is important to consider the specific nature of the generation tasks involved. CODA19 comprises medical articles that utilize specialized medical terminology. Given that most annotators were unfamiliar with this vocabulary, native English speakers did not have a specific advantage over non-native speakers. Additionally, research articles are commonly written in English by authors from various backgrounds. Therefore, it might be that this specific task is robust against the native versus non-native preference. Finally, for the paraphrasing task, we found that native and non-native English annotators handled this task differently. While native English speakers often paraphrased the sentences more freely, non-native English speakers closely followed the original structure, mainly providing synonyms and maintaining the original sentence’s format. An example is added in Appendix J.
@@
Interestingly, only for GPT3.5 adding correct information about the user’s nativeness results in a better performance for the native group versus the non-native group, as shown in Table 3. Qwen even displays the opposite effect, resulting in a lower performance for the native group. The other models are equally preferring both groups. However, when looking into the specific datasets, we again find that the result differs depending on the task. For both GPT3.5 and GPT4o, a better performance is found for native English speakers for StoryCloze and CNN Dailymail, while the other datasets are preferred for non-native speakers. Qwen and Haiku on the other hand, perform better for the non-native English speakers on all datasets, even when given the information that the prompt writer is a non-native speaker. Sonnet performs best for the native group for StoryCloze, all other datasets prefer the non-native group or are on par.
correct information wrong information
model native non-native native non-native
GPT3.5 0.45 0.39 0.39 0.42
GPT4o 0.46 0.46 0.45 0.47
Haiku 0.45 0.45 0.44 0.45
Sonnet 0.45 0.45 0.45 0.45
qwen 7B 0.41 0.42 0.41 0.42
Table 3: Overall average BLEU score of generative tasks per model when adding correct/wrong information about whether or not the prompt was written by a native English speaker.
Additionally, when incorrect information about the user’s nativeness is introduced, we find that all models show improved performance for the non-native group, which is presented to the model as native speakers. This is a remarkable finding, as it emphasizes the importance of the prompt over the user’s nativeness, unlike in classification tasks. GPT3.5 is particularly sensitive to the inclusion of incorrect information regarding the user’s nativeness, with metrics across all datasets increasingly favoring the non-native English speakers and diminishing the performance of the native group when compared to the addition of correct information. GPT4o, still prefers the native group for StoryCloze and CNN Dailymail. The model thus shows similar robustness as for the classification tasks. In this experiment, Haiku and Qwen perform even better for the non-native group. Sonnet, however, is a bit more robust than Haiku.
@@
model guess native guess non-native guess native correctly guess non-native correctly
GPT3.5 0.42 0.42 0.45 0.42
GPT4o 0.46 0.46 0.47 0.46
Haiku 0.44 0.45 0.34 0.43
Sonnet 0.44 0.45 0.39 0.46
Qwen 7B 0.39 0.39 0.43 0.33
Table 4: Overall average BLEU score for all generative tasks per model and group when guessed by the model whether the person is a native or non-native speaker, and when it is guessed correctly.
Table 4 shows the average BLEU scores when the model (correctly) guessed whether or not the writer of the prompt was a native English speaker. Overall, there is no distinct performance variation based on the model’s general guesses. However, when looking at the guesses that were made correctly, differences emerge. All models now perform better for the native group on average, except Haiku and Sonnet. For the individual datasets, we find that CNN Dailymail performs better for the native group, but again CODA19 and Paraphrasing perform better for the non-native group. Similar to the classification tasks, GPT3.5 is more influenced by this additional information than GPT4o.
@@ 6 Discussion
In fact, we find that for the objective classification tasks, the larger models perform better for the native or Western native groups compared to the non-native and not Western native groups respectively. For the subjective classification task, however, we find an opposite effect, where the model interprets prompts of (Western) native English speakers to be more positive than intended. The generative tasks show no clear performance differences in general, but the bias varies per dataset.
Moreover, a strong anchoring effect occurs when the model knows whether or not the prompt writer is a native English speaker, either by correctly guessing, or by inserted information in the system prompt. The bias is so deeply engraved that informing the models of the wrong native and non-native groups, results in a preference towards the group that was indicated as native. The model is largely led by this added information, more than by the prompt itself. However, we find differences between the models. GPT4o appears most resistant to this anchoring effect, while Sonnet on the other hand even changes the language of the response based on this anchor. Furthermore, this anchoring effect seems to be less present for the generative tasks than for the classification tasks.
Interestingly, the results also suggest that the more advanced versions of models, such as GPT4o compared to GPT3.5 and Sonnet compared to Haiku, tend to be more sensitive to the subtleties of native English. On the other hand, we also notice that they are less sensitive to added explicit information of the nativeness or non-nativeness of the annotator, and thus seem to be better hedged against explicitly added bias.
@@ 7 Conclusion
In this work, we analyze bias in LLMs towards native English speakers. More specifically, we analyze whether models perform better for native compared to non-native English speakers. We also analyze whether the models are even further tuned towards Western native English speakers. We find that there are performance differences between native and non-native prompts. More specifically, models become more inaccurate for the latter and this effect is even more pronounced for the Western native English versus not Western native English comparison. Furthermore, we find a strong anchoring effect when information about the user’s nativeness is added or inferred by the model. For our experiments, we used a newly collected dataset consisting of over 12,000 unique prompts from a diverse set of annotators.
@@ 8 Limitations
Our dataset contained a very diverse set of annotators. Nevertheless, it would be interesting to have more study participants for every sub-population, such that general findings at sub-population level could be made as well. Furthermore, our experiments contained mostly annotators having a self-reported level of English of C1 and C2. It would be very interesting to analyze the effects on the performance of LLMs when prompted by people having different levels of English as this will probably also be impactful. Additionally, our results were only gathered for five different models. It would be insightful to extend this analysis to more models, as every model is trained differently and therefore these design choices might lead to different biases within the model. Finally, the main limitation of using LLMs and especially the closed-source variant thereof, is the lack of reproducibility of the results.
@@ Appendix B Annotation set-up
We have set up an annotation platform to gather the annotations. The annotators first get information about the task. They will get a task definition, prompt where part of the answer is marked out with the placeholder [YOUR PROMPT], and the desired output of the LLM. The annotators should complete the prompt such that the desired output would be generated by the LLMs. Figure 6 shows a screenshot of the landing page of the annotation platform together with annotation instructions. An example of an annotation that had to be annotated is shown in Figure 7. We have anonymized all annotations by only providing the self-reported linguistic information in the dataset along with the user id number.
@@ Appendix D Dataset Statistics -Annotations
Set- ids Native or not Western native or not Total
@@ Appendix G Evaluation metrics
BART-score.121212github.com/neulab/BARTScore This evaluation measure approaches the evaluation as a text-generation task, including all parameters learned during the pre-training phase (Yuan et al., 2021).
@@ Appendix H Distribution Amazon food reviews
Figures 8 and 9 show an overview of the wrong predictions of the AmazonFood review dataset for the different groups and models. This shows the distribution between what was predicted and what should be predicted. We only consider here the cases where the model predicted one of the given ratings, and excluded cases where no prediction was given. As shown, for both the native and Western native group, we find a large amount of misclassification for the highest rating. Additionally, neutral is not often predicted for these classes compared to the other groups.
@@ Appendix I Results Sonnet different languages
model amount guess native correctly amount guess non-native correctly

### v2

@@ 5.2 Generative tasks
native non-native western native not western native
GPT3.5 $0.44\pm 0.00$ $0.44\pm 0.00$ $0.44\pm 0.00$ $0.44\pm 0.00$
GPT4o $0.47\pm 0.00$ $0.47\pm 0.00$ $0.47\pm 0.00$ $0.47\pm 0.00$
Haiku $0.45\pm 0.00$ $0.46\pm 0.01$ $0.45\pm 0.01$ $0.46\pm 0.01$
Sonnet $0.45\pm 0.00$ $0.43\pm 0.02$ $0.45\pm 0.00$ $0.45\pm 0.00$
Qwen 7B $0.41\pm 0.00$ $0.44\pm 0.02$ $0.41\pm 0.01$ $0.42\pm 0.01$
Table 4: Overall average BLEU score of the generative task per model and group averaged over three runs and including the standard deviation over these different runs. Significant differences with $p<0.05$ for a Wilcoxon signed-rank test are indicated in bold.
Table 4 shows the average BLEU scores and standard deviation over the three different runs per model and group. The statistical analysis includes all performance metrics. These metrics are included in Appendix L. For the generative tasks, all models perform overall significantly better for the non-native group, and all models except for GPT4o and Sonnet prefer the not western native group.
Despite this clear preference for the non-native group, we do find that the performance highly depends on the specific task at hand. In general, the generative tasks can be divided into two groups: group one with CODA19, which includes a medical research article generation task, and the paraphrasing dataset, group two with StoryCloze and CNN Dailymail. We find that all models prefer the non-native English speakers or not Western native English speakers for group one. Most models prefer (Western) native English speakers for group two. Detailed results are shown in Table 19 in Appendix.
In interpreting these results, it is important to consider the specific nature of the generation tasks involved. CODA19 comprises medical articles that utilize specialized medical terminology. Given that most annotators were unfamiliar with this vocabulary, native English speakers did not have a specific advantage over non-native speakers. Additionally, research articles are commonly written in English by authors from various backgrounds. Therefore, it might be that this specific task is robust against the native versus non-native preference. Finally, for the paraphrasing task, through manual error analysis, we found that native and non-native English annotators handled this task differently. While native English speakers often paraphrased the sentences more freely, non-native English speakers closely followed the original structure, mainly providing synonyms and maintaining the original sentence’s format. An example is added in Appendix J.
@@
(non-)native system prompt guess native system prompt
correct information wrong information overall guess guess correctly
model native non-native native non-native native non-native native non-native
GPT3.5 $0.42\pm 0.01$ $0.40\pm 0.01$ $0.40\pm 0.00$ $0.42\pm 0.01$ $0.42\pm 0.00$ $0.42\pm 0.00$ $0.45\pm 0.01$ $0.42\pm 0.00$
GPT4o $0.47\pm 0.01$ $0.45\pm 0.01$ $0.45\pm 0.00$ $0.47\pm 0.00$ $0.46\pm 0.00$ $0.46\pm 0.00$ $0.48\pm 0.01$ $0.45\pm 0.01$
Haiku $0.45\pm 0.01$ $0.45\pm 0.01$ $0.45\pm 0.01$ $0.45\pm 0.01$ $0.44\pm 0.00$ $0.45\pm 0.00$ $0.34\pm 0.03$ $0.43\pm 0.00$
Sonnet $0.45\pm 0.00$ $0.45\pm 0.00$ $0.45\pm 0.00$ $0.45\pm 0.00$ $0.44\pm 0.01$ $0.45\pm 0.00$ $0.38\pm 0.02$ $0.46\pm 0.01$
Qwen 7B $0.41\pm 0.00$ $0.42\pm 0.00$ $0.41\pm 0.01$ $0.42\pm 0.01$ $0.39\pm 0.01$ $0.39\pm 0.00$ $0.43\pm 0.02$ $0.32\pm 0.01$
Table 5: Overall average BLEU score of generative tasks per model when adding correct/wrong information about whether or not the prompt was written by a native English speaker and when (correctly)guessed by the model whether the person is a native or non-native speaker. We indicated significant differences ($p<0.05$) in bold.
Interestingly, Table 5 shows that adding the correct information about the user’s nativeness affects the GPT-models largely. Now, no significant difference is found between both groups, whereas previously both models significantly preferred the non-native group. However, the result again depends on the specific task. For both GPT3.5 and GPT4o, a better performance is found for native English speakers for StoryCloze and CNN Dailymail, while the other datasets are preferred for non-native speakers. However, this difference between the native and non-native scores has increased for GPT3.5, again showing the anchoring effect. Similar to the classification tasks, this is less visible for GPT4o. While Haiku and Sonnet follow this preference for the native group on these datasets, Qwen always prefers the native group.
Additionally, when incorrect information about the user’s nativeness is introduced, all models show improved or on-par performance for the non-native group, which is presented to the model as native speakers, compared to when they are represented as non-native speakers. This is a remarkable finding, as it emphasizes the importance of the prompt over the user’s nativeness, unlike the results for the subjective classification tasks. The GPT-models are most affected by this as shown in the table.
@@
Table 5 includes the average BLEU scores when the model (correctly) guessed the nativeness of the prompt writer. Overall, no significant differences are found for GPT3.5 and Qwen, contrary to the original experiments in Table 4. Similar to the classification tasks, we find that this is probably due to Qwen’s indecisiveness, which adds extra noise and shows the model sensitivity to useless information as in Shi et al. (2023). When looking at the guesses that were made correctly, GPT4o, Haiku, and Qwen perform significantly better for the native group than for the non-native groups. Sonnet, however still prefers the non-native group, and for GPT3.5 no significant differences are found.
@@ 6 Discussion
We find that GPT4o, Sonnet, and Qwen perform significantly better for the native groups for the objective classification tasks. For the subjective classification task, however, we find an opposite effect, where they interpret prompts of (Western) native English speakers more positively. Similarly, the models significantly prefer the non-native group for generative tasks, but the bias varies per dataset.
Moreover, a strong anchoring effect occurs when the model knows whether or not the prompt writer is a native English speaker, either by correctly guessing, or by inserted information in the system prompt. The bias is so deeply engraved that informing the models of the wrong native and non-native groups, results in a preference towards the group that was indicated as native. The model is largely led by this added information, more than by the prompt itself. However, we find differences between the models. GPT4o appears most resistant to this anchoring effect, while Sonnet on the other hand even changes the language of the response based on this anchor. Furthermore, this anchoring effect seems to be less present for the generative tasks than for the classification tasks. For Qwen, this anchoring effect is a result of the model’s sensitivity to useless information, as this noise results in worse predictions for both groups.
Interestingly, the results also suggest that the more advanced versions of models, such as GPT4o compared to GPT3.5 and Sonnet compared to Haiku, tend to be more sensitive to the subtleties of native English. On the other hand, we also notice that GPT4o is less sensitive to added explicit information of the nativeness or non-nativeness of the annotator, and thus seems to be better hedged against explicitly added bias.
@@ 7 Conclusion
In this work, we analyze bias in LLMs towards native English speakers. We analyze whether models perform better for native compared to non-native English speakers and whether the models are even further tuned towards Western native English speakers. We find that there are significant performance differences between native and non-native prompts. More specifically, models become more inaccurate for the latter and this effect is even more pronounced for the Western native English versus not Western native English comparison for objective classification tasks. Furthermore, we find a strong anchoring effect when information about the user’s nativeness is added or inferred by the model. For our experiments, we used a newly collected dataset consisting of over 12,000 unique prompts from a diverse set of annotators.
@@ 8 Limitations
Our dataset contained a very diverse set of annotators. Nevertheless, it would be interesting to have more study participants for every sub-population, such that general findings at sub-population level could be made as well. Furthermore, our experiments contained mostly annotators having a self-reported level of English of C1 and C2. It would be very interesting to analyze the effects on the performance of LLMs when prompted by people having different levels of English as this will probably also be impactful. Additionally, our results were only gathered for five different models. It would be insightful to extend this analysis to more models, as every model is trained differently and therefore these design choices might lead to different biases within the model. An important limitation of using LLMs and especially the closed-source variant thereof, is the lack of reproducibility of the results. We make available a multilingual dataset, however, have only analyzed the English answers. We leave the analysis of bias in the multilingual dataset for future research. Finally, we acknowledge that our way of evaluating, using the LLM-generated answers as reference sentences for the generative sentences, might be affected by the prompt of the original prompt writer. We have also analyzed the performance when letting the model first annotate the example and then use this to gather the response from the chatbot. However, we find that this resulted in large performance drops for the smaller models. Moreover, for the models that did offer adequate results, we find similar results as when we used the original prompt. The results are included in Appendix N.
@@ 9 Ethical considerations
Acknowledgments
This research was funded by the Statistics Flanders research cooperation agreement on Data Science for Official Statistics.
@@ Appendix B Annotation set-up
We have set up an annotation platform to gather the annotations. The annotators first get information about the task. They will get a task definition, a prompt where part of the answer is marked out with the placeholder [YOUR PROMPT], and the desired output of the LLM. The annotators should complete the prompt such that the desired output would be generated by the LLMs. Figure 3 shows a screenshot of the landing page of the annotation platform together with annotation instructions. An example of an annotation that had to be annotated is shown in Figure 4. An example of the different [Annotator PROMPT] per dataset is shown in Table 7. We have anonymized all annotations by only providing the self-reported linguistic information in the dataset along with the user ID number.
@@
Dataset Example Prompt
Abductivenli Beginning: Mike was in the car on the highway. Middle 1 : [Annotator Prompt]. Middle 2: A deer never ran in front of his car. Ending: When he got to the hospital, they saw that it had been broken
Story Cloze Sentence1: [Annotator Prompt] Sentence2: Suddenly, there was an announcement. Sentence3: The school was on a lockdown. Sentence4: The kids sat quietly, and waited.
Timetravel Sentence 1: Little Charlie and his dad were painting the garage. Sentence 3: His dad turned around and started to laugh Sentence 4: Charlie had paint on him from head to toe Sentence 5: His dad rinsed him off with water from the hose Option 1: [Annotator Prompt] Option 2: Charlie had some trouble controlling the brush.
AmazonFood rating This is [Annotator Prompt]
McTaco Sentence: The legitimization of gambling led to its increased legalization across the US. Question: [Annotator Prompt]
TweetQA Context: Praying for everyone here in Vegas. I witnessed the most unimaginable event tonight. We are okay. Others arent. Please pray. –Jake Owen (@jakeowen) October 2, 2017 Question: [Annotator Prompt] Answer: people were not okay
CNN DailyMail [Annotator Prompt]
CODA19 [Annotator Prompt]
Paraphrase [Annotator Prompt]
Commonsense I walk under the park. [Annotator Prompt]
Table 7: Example of a prompt to annotate per dataset. [Annotator Prompt] indicates where the prompt of that the annotator should come up with, should fit in the text.
@@ Appendix D Dataset Statistics -Annotations
Set ids Native or not Western native or not Total
@@ Appendix G Evaluation metrics
BART-score.131313github.com/neulab/BARTScore This evaluation measure approaches the evaluation as a text-generation task, including all parameters learned during the pre-training phase (Yuan et al., 2021).
@@ Appendix H Distribution Amazon food reviews
Figures 5 and 6 show an overview of the wrong predictions of the AmazonFood review dataset for the different groups and models for one of the three runs. This shows the distribution between what was predicted and what should be predicted. We only consider here the cases where the model predicted one of the given ratings, and excluded cases where no prediction was given. As shown, for both the native and Western native group, we find a large amount of misclassification for the highest rating. Additionally, neutral is not often predicted for these classes compared to the other groups.
@@ Appendix I Results Sonnet different languages
amount guess native correctly amount guess non-native correctly